A circle has center (4, 2). Prove that (0,3) and (3, -1) are not both points on the circle.
Answer #1

Problem in answer then comment below...i will help you..

.

Please thumbs up for this solution..thanks..

.

center = (4,2) ..

first point is (0,3) .. distance between center and this point is sqrt (16+1)= sqrt(17)

2nd point is (3,-1) .. distance between center and this point is sqrt (1+9)= sqrt(10) ..

here these two distances not same .. but if these two points are on circle then these distances should be same because in that case this distance is radius .. so

these points are not both on the circle ..